Robin ran 5/1/3 miles on Sunday, 2/1/4 miles on Tuesday, and 3/1/2 miles on Friday. How far did she run during the week?: Robin ran 5/1/3 miles on Sunday, 2/1/4 miles on Tuesday, and 3/1/2 miles on Friday. How far did she run during the week?
Answer by vleith(1174)

You need to get the daily mileage over a common denominator to add them.
You can do that by putting the full amount from each day in fractional form, or just adding the integer parts and then adding the fractional parts. Let's add the two parts separately
(5  + 1/3 ) + ( 2 + 1/4) + (3 + 1/2)
(5+2+3) + (1/3 + 1/4 + 1/2)
 10 + (4/12 + 3/12 + 6/12)
 10 + (13/12)
 10 + ( 12/12 + 1/12)
 10 + 1 + 1/12
11 and 1/12 miles
Question 160787: Robin ran 5/1/3 miles on Sunday, 2/1/4 miles on Tuesday, and 3/1/2 miles on Friday. How far did she run during the week?: Robin ran 5/1/3 miles on Sunday, 2/1/4 miles on Tuesday, and 3/1/2 miles on Friday. How far did she run during the week?
Answer by KnightOwlTutor(216)

Convert mixed numbers into fractions:
5&1/3=16/3
2&1/4=9/4
3&1/2=7/2
Let's determine what the least common denominator would be. Upon visual inspection I pick 12.
Make the appropriate modifications to the fraction without changing the ratio
(16/3)(4/4)=64/12
(9/4)(3/3)=27/12
(7/2)(6/6)=42/12

Now add each fraction together to get the # of miles run during the week
(64+27+42)/12=11 and 1/12 miles